# EmoSphere++: Emotion-Controllable Zero-Shot Text-to-Speech via Emotion-Adaptive Spherical Vector

### v1

@@ V-E2 Effectiveness of Predicted versus Real VAD Values
To compare the impact of SER models that are trained with emotion dimension annotations, we conducted comparative experiments on the interactive emotional dyadic motion capture database (IEMOCAP) dataset [59], which includes emotional dimension annotations. As a training input for the emotion-adaptive coordinate transformation, we compared two features: real VAD values from human labeled annotations and predicted VAD values from the SER. Table VI shows that the objective metrics of prosodic expressiveness for both real and predicted VAD values are similar, suggesting that the predicted values perform comparably to manual annotations. Furthermore, we demonstrated the pitch tendency plot in Fig. 9, where the predicted VAD values reveal a consistent and precise pattern across all emotions. The results indicate that the VAD values predicted from the SER provide reliable emotion style and intensity modeling comparable to the real VAD values. These results indicate that SER-based VAD extraction offers a dependable alternative to manual annotations, maintaining similar performance while improving control over emotional nuances.
@@
V-E3 Comparison of Coordinate Transformation
To compare effective modeling, we compared the mean-based method of the spherical emotion vector (SEV) [3] and EASV. For a fair comparison, we compared intensity accuracy using speech pairs of the same emotion and style from the same model, differing only in the coordinate transformation. We referred to values of W as weak, M as medium, and S as strong. As shown in Fig. 10, all types of emotion intensity pairs (W$<$M, M$<$S, and W$<$S) demonstrate consistently high average accuracy for individual emotions and overall. These results suggest that, while the mean-based approach of SEV captures the target emotion to a certain extent, it can be further improved by considering the distribution of other emotional categories, such as variance. Our proposed EASV method builds on the SEV, leading to a more comprehensive emotional representation. Furthermore, the distributions of the target and other emotional categories demonstrate the effectiveness of modeling style and intensity for each
@@ VI-A Limitations of Data Imbalance
VI-B Remaining Challenges of SER Model
As we summarized the SER models in Section III-A, we utilized the fine-tuned wav2vec 2.0 for an SER task. In the study [48], VAD predicted by the wav2vec 2.0-based SER model was significant. Hence, we utilized VAD pseudo-labels to avoid the inherent subjectivity and high costs associated with data collection. However, the approach relies on the performance of the trained SER model and similar biases and challenges as those encountered in SER [48, 68]. We expect the extended VAD predictor [69] will solve this problem. Although we only experimented with predicting VAD values here, we can extend this to combine real VAD values across different datasets.
@@ VII Conclusion
In this paper, we presented EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We presented a comprehensive analysis and study that characterizes emotional states as derivatives of primary emotions using a dimensional model at the initial stage. We propose a novel EASV to model emotional style and intensity as derivatives of primary emotions. We also introduced a controllable and expressive speech synthesis framework that utilizes a multilevel style encoder with additional loss functions. The experimental results thoroughly analyze the components of our model and demonstrate its ability to effectively synthesize and control speech performance, even in zero-shot scenarios. We demonstrated that controlling the spherical vector along the VAD axis enables independent emotional style and intensity control. Through ablation studies, we also showed that the proposed EASV effectively synthesizes the complex nature of emotion. This article only focused on studying emotion-controllable TTS for a limited set of emotions. Our proposed spherical vector can enable complete emotion control in most existing emotional speech synthesis frameworks. Furthermore, we plan to expand our experiments to include emotional voice conversion in future studies.

### v2

@@ V-E2 Effectiveness of Predicted versus Real VAD Values
To compare the impact of emotional attribute prediction models [53], we conducted comparative experiments using the IEMOCAP dataset [65], which includes emotion dimension annotations. As a training input for the emotion-adaptive coordinate transformation, we compared two features: real VAD values from human labeled annotations and predicted VAD values from the emotional attribute prediction [53]. Table VI shows that the objective metrics of prosodic expressiveness for both real and predicted VAD values are similar, suggesting that the predicted values perform comparably to manual annotations. Additionally, as shown in Fig. 9, both the real and predicted VAD values exhibit a consistent and accurate pattern across all emotions, aligning well with the ground truth in intensity control. The results indicate that the VAD values predicted from the emotional attribute prediction [53] provide reliable emotion style and intensity modeling comparable to the real VAD values.
V-E3 Comparison of Coordinate Transformation
For a fair comparison, we compared intensity accuracy using speech pairs generated from the same model, with identical emotion and style, differing only in their coordinate transformation approach. As shown in Fig. 10, all types of emotion intensity pairs (W$<$M, M$<$S, W$<$S) demonstrate accuracy across individual emotions, styles, and overall. We referred to values of W as weak, M as medium, and S as strong. This experiment was conducted to evaluate whether more precise modeling of emotion style and intensity leads to clearly distinguishable synthesized speech when given pseudo-labels as input. By assessing the accuracy of intensity differentiation, we aim to validate the effectiveness of EASV in generating perceptually distinct emotional variations. The EASV method demonstrates consistently high average accuracy for individual emotions and overall. These results suggest that, while the mean-based approach of SEV captures the emotion to a certain extent, it remains unstable at specific intensity levels and styles. This indicates that considering the distribution of other emotional categories in EASV further improves the modeling of emotion style and intensity.
@@ VI-A Limitations of Data Imbalance
VI-B Remaining Challenges of Emotional Attribute Prediction Model
As we summarized the emotional attribute prediction models in Section III-A, we utilized the fine-tuned wav2vec 2.0 for an emotional attribute prediction [53] task. In the study, VAD predicted by the wav2vec 2.0-based emotional attribute prediction model [53] was significant. Hence, we utilized VAD pseudo-labels to avoid the inherent subjectivity and high costs associated with data collection. However, the approach relies on the performance of the trained emotional attribute prediction model and contains similar biases and challenges to those encountered in emotional attribute prediction [53]. We expect the extended VAD predictor [24] to mitigate the biases inherent in emotional attribute prediction while providing more accurate VAD estimations, ultimately addressing this issue.
@@ VII Conclusion
In this paper, we presented EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. To achieve this, we proposed the novel emotion-adaptive spherical vector (EASV), which models emotional style and intensity as derivatives of primary emotions. Building on this, our comprehensive analysis validates that a dimensional model can characterize emotional states in relation to primary emotions. Additionally, a zero-shot speech synthesis framework with rich expressiveness and controllability was developed, utilizing a joint attribute style encoder with additional loss functions, without being restricted by predefined speaker and emotion labels. The experimental results thoroughly analyze the components of our model and demonstrate its ability to effectively synthesize and control speech performance, even in zero-shot scenarios. We demonstrated that by controlling the spherical vector along the VAD axis, explicit adjustments to emotional style and intensity enable fine-grained emotional expression. Ablation studies further confirm that the proposed EASV effectively synthesizes the complex nature of emotion. While this article only focused on studying emotion-controllable TTS for a limited set of emotions, our proposed spherical vector can enable complete emotion control in most existing emotional speech synthesis frameworks. Future work will expand these experiments to include emotional voice conversion.